If 8.28 L of hydrogen gas at SATP is produced, how many moles of aluminium were reacted?


1
Expert's answer
2022-06-13T02:53:04-0400

under normal conditions occupies a volume of 22.4 l per 1 mol gas. so we find how many moles are in 8.28 l of H2 :

n (H2) = V / 22,4 = 8,28 / 22,4 = 0,37 mol


2Al + 6HCl = 2AlCl3 + 3H2

2mol(54 g) Al —> 3mol H2

X g Al —> 0,37 mol H2


X = 54 * 0,37 / 3 = 6,66 g



Answer: 6,66 g Al